# OpenLVLM-MIA: A Controlled Benchmark Revealing the Limits of Membership Inference Attacks on Large Vision-Language Models

### v1

@@ Abstract
OpenLVLM-MIA is a new benchmark that highlights fundamental challenges in evaluating membership inference attacks (MIA) against large vision-language models (LVLMs). While prior work has reported high attack success rates, our analysis suggests that these results often arise from detecting distributional bias introduced during dataset construction rather than from identifying true membership status. To address this issue, we introduce a controlled benchmark of 6,000 images where the distributions of member and non-member samples are carefully balanced, and ground-truth membership labels are provided across three distinct training stages. Experiments using OpenLVLM-MIA demonstrated that the performance of state-of-the-art MIA methods converged to random chance under unbiased conditions. By offering a transparent and unbiased benchmark, OpenLVLM-MIA clarifies the current limitations of MIA research on LVLMs and provides a solid foundation for developing stronger privacy-preserving techniques.
@@ 1.1 Escalating Privacy Risks
Training LVLMs requires large amounts of image data; thus, large-scale datasets such as LAION-5B [24] and Conceptual Captions [25] are used for training. Since these datasets are collected via web crawling, there is a risk that confidential images such as medical images, personal photos, and copyrighted works may be unintentionally included in the training data. Private data like patient diagnostic images and passport photos, as well as intellectual property such as artworks and commercial photographs, may be used without the owners’ consent.
@@ 1.2 Limitations of Existing MIA Research
In recent years, several MIA techniques for LLMs [26, 6] have been proposed, reporting high success rates. For LVLMs as well, Li et al. [13] achieved a maximum AUROC of 0.82 on their VL-MIA benchmark, which suggested that LVLMs present substantial privacy risks.
However, our analysis revealed that these high performance results may not necessarily reflect true membership detection.
@@
2) Uncertain ground truth: Because many LVLMs keep parts of their training data undisclosed, the true membership of test data cannot be verified.
@@ 1.3 Motivation and Goal
The above two observations raise a research question: What do current MIA methods actually measure? Does a high attack success rate truly detect model memory, or is it merely capturing biases in the data during dataset construction? To answer the research question, we construct a new controlled benchmark dataset - OpenLVLM-MIA.
@@ 1.4 Contributions
1) OpenLVLM-MIA benchmark: We introduced a controlled dataset comprising 6,000 images characterized by three defining properties: (i) the member and non-member subsets are distributionally matched; (ii) each example is annotated with a ground-truth membership label; and (iii) the dataset supports independent evaluation at multiple training stages. This design reduces distributional bias, enabling fair evaluation of membership inference attacks (MIAs).
@@ 2.1 Membership Inference Attacks
2.2 MIA Benchmarks for Vision–Language Models
@@ 2.2.1 VL-MIA Dataset
VL-MIA [13] was introduced as the first MIA benchmark for LVLM, which targets the following three LVLM systems:
  • LLaVA v1.5 [15]
  • MiniGPT-4 [32]
  • LLaMA-Adapter v2 [7]
VL-MIA consists of two subsets:
@@
Since the training data for these models is not fully public, non-members must be drawn from temporally mismatched or distributionally different sources (e.g., DALL-E–generated images). Consequently, MIA methods may detect temporal shifts or differences between generated and natural images rather than true membership.
@@ 3.2 Distribution Bias Problem
MIA evaluation faces a challenge called dataset distribution bias, which arises when the member and non-member samples are drawn from different distributions, such as the followings[3]:
@@ 4 Proposed Benchmark Model: OpenCLIP-LLaVA
4.1 Limitations of Prior Work and Mitigations in the Proposed Benchmark
@@ 4.3 Public Resources
We will ensure reproducibility by releasing the following:
@@ 5.1 Dataset Design
The proposed OpenLVLM-MIA benchmark enables independent evaluation of membership across the entire training pipeline, i.e., vision encoder pretraining, projector pretraining, and instruction tuning, because LVLM training typically proceeds in multiple stages, each relying on a different dataset. The key features are as follows:
@@ 6.1.3 Results
Figure 1 visualizes the embedding space in 3D. Nonlinear dimension reduction methods such as t-SNE and UMAP emphasize local structure, but image embeddings are high-dimensional, making it difficult to visualize separation in 3D space using unsupervised methods. Therefore, in this study, we quantified separability using supervised dimensionality reduction with membership labels, where the Fisher discriminant axis $w\propto S_{w}^{-1}(\mu_{1}-\mu_{0})$ derived from LDA [5] serves as the first axis. Furthermore, by applying PCA [21] in the residual space orthogonal to this, we constructed the second and third axes, thereby visualizing variations beyond the discriminant axis. This makes any separation (or its absence) visible.
Figure 1: LDA+PCA 3D visualization of the embedding space. dim1 is the supervised LDA discriminant axis; dim2–3 are the top two PCA components in the subspace orthogonal to the LDA axis. VL-MIA shows clear separation, while OpenLVLM-MIA exhibits substantial overlap.
@@ 6.2.2 Results
In our carefully controlled benchmark, all state-of-the-art MIA methods converge to AUROC 0.41–0.52, demonstrating performance indistinguishable from random guessing (Table 4). The best case is 0.527 (Max Rényi 0%, $\alpha{=}0.5$), while the worst is 0.407 (Min-0%, Projector, inst). These results suggest that, in this setting, membership inference is effectively uninformative.
@@
In the Vision Encoder phase, AUROC ranges from 0.450 to 0.526; Max Rényi methods are relatively higher but remain around 0.52. The Projector phase shows the lowest performance overall, with AUROC 0.407–0.513 (minimum 0.407 for Min-0% on inst); TPR@0.05FPR is below 0.04 in over half the conditions. In the Instruction Tuning phase, AUROC narrows to 0.484–0.527 and differences between methods are minimal.
@@ 7.3 Analysis by Training Phase
Projector stage: This stage shows the lowest AUROC (0.407–0.513). Although the dataset used here showed the largest distribution bias in the OpenLVLM-MIA audit, this stage’s contribution to membership signals appears small. The effects of multimodal integrated learning and the fact that the weight of the projector is small relative to the entire model, resulting in a low proportion of training data interfering with the model, can also be considered contributing factors.
@@ 7.4 Future Work
Current MIA approaches for large-scale foundational models are primarily limited to pure applications designed for language models and do not account for multimodality. Future development requires methods that explicitly leverage the interaction between vision and language. Analyzing patterns in caption generation for specific images and the consistency between visual attributes and linguistic descriptions may enable more effective membership inference. In addition, establishing standardized benchmarks with minimal distributional bias, such as OpenLVLM-MIA, is essential.
@@ 8 Limitations
The membership inference attacks (MIA) we evaluate assume a gray-box setting with access only to output probabilities; white-box attacks are out of scope.

### v2

@@ Abstract
OpenLVLM-MIA is a new benchmark that highlights fundamental challenges in evaluating membership inference attacks (MIA) against large vision-language models (LVLMs). While prior work has reported high attack success rates, our analysis suggests that these results often arise from detecting distributional bias introduced during dataset construction rather than from identifying true membership status. To address this issue, we introduce a controlled benchmark of 6,000 images where the distributions of member and non-member samples are carefully balanced, and ground-truth membership labels are provided across three distinct training stages. Experiments using OpenLVLM-MIA demonstrated that the performance of state-of-the-art MIA methods approached chance-level. OpenLVLM-MIA, designed to be transparent and unbiased benchmark, clarifies certain limitations of MIA research on LVLMs and provides a solid foundation for developing stronger privacy-preserving techniques.
@@ 1.1 Escalating Privacy Risks
Training LVLMs requires large amounts of image data; thus, large-scale datasets such as LAION-5B [25] and Conceptual Captions [26] are used for training. Since these datasets are collected via web crawling, there is a risk that confidential images, such as medical images, personal photos, and copyrighted works, may be unintentionally included in the training data. Private data like patient diagnostic images and passport photos, as well as intellectual property, such as artworks and commercial photographs, may be used without the owners’ consent.
@@ 1.2 Limitations of Existing MIA Research
In recent years, several MIA techniques for LLMs [27, 7, 4] have been proposed, reporting high success rates. For LVLMs as well, Li et al. [14] achieved a maximum AUROC of 0.82 on their VL-MIA benchmark, which suggested that LVLMs present substantial privacy risks. However, our analysis suggests that these high performance results may not necessarily reflect true membership detection.
@@
2) Uncertain ground truth: Because many LVLMs keep parts of their training data undisclosed, the true membership of test data cannot be verified, making evaluation based on such benchmarks potentially ambiguous.
@@ 1.3 Motivation and Goal
The above two observations raise a research question: What do current MIA methods actually measure? Does a high attack success rate truly detect model memory, or is it merely capturing biases in the data during dataset construction? To answer the research question, we construct OpenLVLM-MIA, a new controlled benchmark consisting of a benchmark dataset and a target model.
@@ 1.4 Contributions
1) OpenLVLM-MIA benchmark dataset: We introduced a controlled dataset comprising 6,000 images characterized by three defining properties: (i) the member and non-member subsets are distributionally matched; (ii) each example is annotated with a ground-truth membership label; and (iii) the dataset supports independent evaluation at multiple training stages. This design reduces distributional bias, enabling fair evaluation of membership inference attacks (MIAs).
@@ 2.1 Membership Inference Attacks
2.2 MIA Benchmark Datasets for Vision–Language Models
@@ 2.2.1 VL-MIA Dataset
VL-MIA [14] is the first MIA benchmark for LVLMs, targeting three systems (LLaVA v1.5 [16], MiniGPT-4 [34], and LLaMA-Adapter v2 [8]). It contains 600 member and 600 non-member samples split into two subsets:
@@
Since the training data for the above three systems is not fully public, non-members must be drawn from temporally mismatched or distributionally different sources (e.g., DALL-E–generated images). Consequently, MIA methods may detect temporal shifts or differences between generated and natural images rather than true membership.
@@ 3.2 Distribution Bias Problem
MIA evaluation faces a challenge called dataset distribution bias, which arises when the member and non-member samples are drawn from different distributions. In this paper, we use the term temporal shift as distribution differences caused by the time at which data were collected, which can lead to differences in the appearance and characteristics of the images.
Examples include:
@@ 4 Proposed Benchmark Model: OpenCLIP-LLaVA
4.1 Limitations of Prior Work and Mitigations in the Proposed Benchmark Model
@@ 4.2.2 Performance comparison on LVLM Benchmarks
Scores are computed using the official evaluation protocol/scripts of each benchmark (higher is better, $\uparrow$).
@@ 4.3 Public Resources
We ensure reproducibility by releasing the following resources. Note that the dataset and model weights are hosted on Hugging Face and can be accessed using the code provided in the GitHub repository.
  • Evaluation scripts and implementation code
    https://github.com/yamanalab/openlvlm-mia
  • The OpenLVLM-MIA dataset and OpenCLIP-LLaVA (7B) model weights
    Accessible via the code provided in the repository above.
@@ 5.1 Dataset Design
The proposed OpenLVLM-MIA benchmark dataset enables independent evaluation of membership across the entire training pipeline, i.e., vision encoder pretraining, projector pretraining, and instruction tuning, because LVLM training typically proceeds in multiple stages, each relying on a different dataset. The key features are as follows:
@@ 6.1.3 Results
Figure 1 visualizes the embedding space in 3D. Nonlinear dimension reduction methods such as t-SNE and UMAP emphasize local structure, but image embeddings are high-dimensional, making it difficult to visualize separation in 3D space using unsupervised methods. Therefore, in this study, we quantified separability using supervised dimensionality reduction with membership labels, where the Fisher discriminant axis $w\propto S_{w}^{-1}(\mu_{1}-\mu_{0})$ derived from LDA [6] serves as the first axis. Furthermore, by applying PCA [22] in the residual space orthogonal to this, we constructed the second and third axes, thereby visualizing variations beyond the discriminant axis. This makes any separation (or its absence) visible. The plots visually support the findings: VL-MIA displays a noticeable separation between member and non-member samples, whereas OpenLVLM-MIA exhibits considerable overlap.
Figure 1: LDA+PCA 3D visualization of the embedding space. dim1 is the supervised LDA discriminant axis; dim2–3 are the top two PCA components in the subspace orthogonal to the LDA axis. VL-MIA shows noticeable separation, while OpenLVLM-MIA exhibits considerable overlap.
@@ 6.2.2 Results
In our carefully controlled benchmark, all state-of-the-art MIA methods converge to AUROC 0.41–0.52, demonstrating performance indistinguishable from random guessing (Table 4). The best case is 0.527 (Max Rényi 0%, $\alpha{=}0.5$), while the worst is 0.407 (Min-0%, Projector, inst). These results suggest that, in this setting, membership inference is effectively uninformative under our controlled benchmark.
@@
In the Vision Encoder phase, AUROC ranges from 0.450 to 0.526; Max Rényi methods are relatively higher but remain around 0.52. The Projector phase shows the lowest performance, with AUROC 0.407–0.513 (minimum 0.407 for Min-0% on inst); TPR@0.05FPR is below 0.04 in over half the conditions. In the Instruction Tuning phase, AUROC narrows to 0.484–0.527 and differences between methods are minimal.
@@ 7.3 Analysis by Training Phase
Projector stage: This stage shows the lowest AUROC (0.407–0.513). Although the dataset used here showed the largest distribution bias in the audit, this stage’s contribution to membership signals appears small. The effects of multimodal integrated learning and the fact that the weight of the projector is small relative to the entire model, resulting in a low proportion of training data interfering with the model, can also be considered contributing factors.
@@ 7.4 Future Work
Current MIA approaches for large vision-language models are primarily limited to pure applications designed for language models and do not account for multimodality. Future development requires methods that explicitly leverage the interaction between vision and language. For example, AMIA [32] uses text-image correlation for defense. Inverting such mechanisms could enable stronger attacks by detecting sensitivity in specific image regions or susceptibility to prompt variations. Analyzing patterns in caption generation under varying instructions and the consistency between visual attributes and linguistic descriptions may enable more effective membership inference. In addition, establishing standardized benchmarks with minimal distributional bias, such as OpenLVLM-MIA, is essential.
@@ 8 Limitations
The membership inference attacks (MIA) we evaluate assume a gray-box setting using standard prompts, focusing on output probabilities rather than prompt engineering [16]; white-box attacks are out of scope.